1.How many quarts of pure alcohol must be added to 10 quarts of a 15% alcohol solution to obtain a mixture which is 25% alcohol? All percentages are by volume: 1.How many quarts of pure alcohol must be added to 10 quarts of a 15% alcohol solution to obtain a mixture which is 25% alcohol? All percentages are by volume
Answer by ankor@dixie-net.com(4538)

1.How many quarts of pure alcohol must be added to 10 quarts of a 15% alcohol solution to obtain a mixture which is 25% alcohol? All percentages are by volume.
:
Let x = amt of pure alcohol required
:
Write amt of alcohol equation:
.15(10) + x = .25(x+10)
:
1.5 + x = .25x + 2.5
:
x - .25x = 2.5 - 1.5
:
.75x = 1
x = 1/.75
x = 11/3 qt of pure alcohol required
;
:
Check solution on cal
1.5 + 1.33 = .25(2.83)
2.83 = 2.83